OpenStudy (anonymous):

A coin is biased such that it results in 2 heads out of every 3 coin flips, on average. Which statement is true? a)The theoretical probability of getting heads from an unbiased coin is about 66%. b)The sequence HTHTTHTTT is valid for the biased coin. c)The biased coin has about a 33% probability of displaying tails. d)The sequence HTTHTHTHT is valid for the biased coin.

OpenStudy (kirbykirby):

Do you have any ideas?

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

i think

OpenStudy (kirbykirby):

yeah :)

OpenStudy (anonymous):

yaaaay!

OpenStudy (perl):

P(H) = 2/3 , P(T) = 1/3

OpenStudy (perl):

so c)